# LDI: Localized Data Imputation

### v1

@@ Abstract
Missing values are a common challenge in real-world tabular data and can significantly impair downstream analysis. While Large Language Models (LLMs) have recently shown promise in data imputation, existing methods often rely on broad, unfiltered prompts that compromise accuracy, scalability, and explainability. We introduce LDI (Localized Data Imputation), a novel framework that improves both the accuracy and transparency of LLM-based imputation by selecting a compact, contextually relevant subset of attributes and tuples for each missing value. This localized prompting reduces noise, enables traceability by revealing which data influenced each prediction, and is effective across both hosted LLMs and lightweight local models. Our extensive experiments on four real-world datasets show that LDI outperforms state-of-the-art methods, achieving up to 8% higher accuracy when using hosted LLMs. The gains are more substantial with lightweight local models, reaching nearly 17% and 97% accuracy on some datasets when using 3 and 10 examples, respectively. In addition to higher accuracy, LDI offers improved interpretability and robustness to data inconsistencies, making it well-suited for high-stakes and privacy-sensitive applications.
Index Terms:
Data Imputation, Dependency Evaluation, Explainability, Data Wrangling
I Introduction
@@
Missing values are a common characteristic of many real-world datasets, particularly those derived from open or heterogeneous sources. If not properly addressed, they can significantly undermine the reliability of downstream analysis and machine learning applications. For example, many learning algorithms operate under the assumption that all input data is fully observed [1, 2, 3], and their performance often degrades when faced with incomplete records. This not only reduces predictive accuracy but can also introduce systematic bias into analytical outcomes [4, 5, 6, 7]. As a result, predicting or imputing missing values is a critical component of the broader data wrangling process, often helping to preserve the underlying structure of the data and ultimately leading to improved performance in downstream tasks [8].
@@
Example 1 (Motivating Example): Consider a restaurant table with missing values in the City column, as shown in Figure 1. There are several other attributes—including Description, Phone, and Reviews—that can potentially help infer the correct city for each missing entry, but it is not clear which of these attributes are actually informative or reliable for this purpose. The Description column may contain location-related hints—e.g., “Modern American dishes” and “Elegant Italian dining”—but these are typically too vague to pinpoint a specific city. In contrast, the Reviews field may sometimes offer more concrete clues, such as “Best gelato in Vegas”, though such references are often embedded in noisy or irrelevant text. The Phone column, especially the area code, can also provide strong signals about geographic location. However, inconsistencies in phone number formatting make it difficult to reliably extract and compare these patterns across entries.
Figure 1: Example of data imputation
@@
Existing approaches to data imputation range from simple heuristics to advanced generative models. Heuristic techniques—such as mean, median, or mode imputation—are fast and easy to implement [9], but they treat attributes independently and overlook complex relationships, often leading to biased or inaccurate results [10]. Model-based methods, including regression, expectation-maximization (EM), and matrix factorization, attempt to infer missing values by modeling statistical dependencies among attributes [11]. While generally more accurate, they can be computationally expensive and rely on strong assumptions about the data distribution. Recent deep learning–based approaches, such as autoencoders, GANs, and transformer models, have shown promise in capturing nonlinear patterns, especially in large or multi-modal datasets [12]. Building on this progress, large language models (LLMs) have emerged as powerful tools for data imputation, thanks to their ability to model complex dependencies and generalize across diverse contexts [13, 14, 15].
However, directly applying LLMs to missing value imputation presents several challenges. First, the full content of a database table often exceeds the model’s context window, making it infeasible to include all relevant information. Second, even when the context window is sufficient, providing the entire table can introduce substantial noise, overwhelming the model with irrelevant or weakly related details, potentially degrading performance. Third, in large tables, it becomes difficult to determine which parts of the input influenced the imputation, limiting the transparency and explainability of the process. These challenges raise a key research question: Can we identify and extract a minimal yet sufficient subset of the table that enables more accurate and interpretable imputations? On what factors does this selection depend, and how can we ensure that the extracted context preserves the necessary statistical or semantic cues for reliable inference?
@@
We hypothesize that, in many cases, missing values can be imputed accurately using a localized context—that is, a small subset of the table that contains the most relevant information for the missing entry. Our approach is to identify this localized context prior to imputation by decomposing the problem into two sub-tasks: (1) selecting a subset of columns that are most relevant to the column with missing values, and (2) identifying a subset of rows that provide sufficient contextual evidence for imputation. Our extensive evaluation across a variety of datasets demonstrates that using this targeted strategy not only improves performance but also helps explain the predictions by showing which parts of the data influenced the output.
Contributions: Our contributions can be summarized as follows: (1) We propose a method for selecting a small, informative subset of attributes and tuples to guide LLM-based imputation, improving both accuracy and explainability. (2) Our subset selection is designed to handle inconsistencies in real-world data during the selection process, making it more robust to noise. (3) We present a complexity analysis of the proposed method, demonstrating its suitability for large-scale datasets. (4) We show that our method can be effectively applied to small local LLMs, offering a practical solution for privacy-sensitive applications. (5) Through extensive experiments on real-world datasets from various domains, we demonstrate that LDI outperforms existing state-of-the-art approaches in terms of accuracy and offers higher explainability compared to other LLM-based methods. (6) To support reproducibility and further research, we publicly release our code, framework, and datasets. [^1] https://github.com/soroushomidvar/LDI
II Related Works
@@ II-A Data Imputation
Traditional Methods: Basic approaches replace missing values with fixed constants, such as zero or the mean, or use strategies like Last Observation Carried Forward (LOCF) and Next Observation Carried Backward (NOCB) [16, 17, 18]. These approaches are computationally inexpensive and straightforward to implement but rely on restrictive assumptions, such as the randomness of missing values or ordering of tuples, that often lead to biased results [19], limiting their applicability in real-world scenarios where data is complex and irregular.
@@
Statistical Models: While basic methods rely on direct substitution, statistical models enhance imputation accuracy by estimating the underlying data distribution and predicting missing values accordingly. Examples include approaches based on relational dependency networks [20], factor graphs [21, 22], distance likelihood maximization [8], low-rank matrix completion [23], and non-linear models such as tree-based ensembles [24, 25]. While effective in certain scenarios, these approaches often rely on assumptions about data properties, such as normality or linearity, which can limit their applicability to real-world datasets.
@@
similarity-based Approaches: Another line of research focuses on similarity-based techniques, where missing values are inferred based on similar neighbors with complete values. Neighbor-based methods, such as k-Nearest Neighbors (KNN) [26] and its extension, KNN Ensemble (KNNE) [27], find the nearest neighbors with the smallest distance on complete attributes and aggregate their values as the imputation for the missing value. Similarly, MIBOS [28] imputes missing values through a similarity-based iterative approach that counts matching attribute values. Instead of relying on individual neighbors, some methods group similar tuples into clusters and perform the imputations within each cluster. For example, CMI [2] imputes missing values by first clustering the dataset and then generating plausible values using similar tuples in the cluster. MGI [29] applies the Gaussian mixture model (GMM), and IFC [30] uses iterative fuzzy clustering to integrate mean imputation, hot deck, and multiple imputation. However, these similarity-based methods are constrained by their similarity functions, cannot handle semantic similarity, and have limited performance if used directly on the data without additional steps.
@@
Generative Models: Beyond these traditional methods, recent advances in generative models have opened new possibilities for data imputation by enhancing the ability to learn data patterns and generate more realistic values. For instance, GAIN [12] uses the Generative Adversarial Nets (GAN) framework, where a generator imputes missing components and a discriminator ensures the generated data follows the true distribution. MIDA [31] proposes a multiple imputation model using deep denoising autoencoders, and HI-VAE [32] adapts variational autoencoders to work with heterogeneous data. IPM [33] leverages pre-trained language models to capture semantics for predicting missing values in databases. It introduces IPM-Multi, which models imputation as multiclass classification, and IPM-Binary, which uses binary classification for candidate selection. A more recent work, FMW [13], also demonstrates that LLMs can handle missing data imputation effectively by framing them as prompting tasks. Although these models achieve state-of-the-art results, their black-box nature lacks explainability, making it difficult to understand how and why specific imputations are generated. The inability to trace an imputed value back to a specific part of the input data prevents validating, correcting, and refining imputations and limits the model’s applicability where understanding why a particular value was imputed can be just as important as the imputed value itself.
II-B Dependency Detection
@@
In the context of attribute selection, our method is closely related to research on identifying dependencies between attributes, particularly through the use of Functional Dependencies (FDs) and Approximate Functional Dependencies (AFDs). FDs describe exact relationships where the value of one attribute is fully determined by another. Formally, an FD $X\rightarrow A$ holds in a relation $r$ if for all tuples $t$ and $u$ in $r$, $t[X]=u[X]$ implies $t[A]=u[A]$. FDs are widely used in database design to capture exact deterministic relationships [34, 35, 36, 37]. In practice, however, FDs are often too strict due to inconsistencies or noise in real-world data. To address this, AFDs have been introduced [38, 39]. An AFD $X\rightarrow A$ holds approximately if it is satisfied by most of the data, with the quality measured by the error $e(X\rightarrow A)$, defined as the minimum fraction of tuples that must be removed to make the dependency exact [40].
@@
FDs and AFDs form the basis of several data cleaning techniques, including sparse regression for dependency discovery [41], conditional functional dependencies [35], and dynamic AFDs that tolerate dirty data by incorporating data source information [42]. However, these approaches are generally designed for structured attributes with consistent formatting and are less effective when applied to unstructured, and possibly lengthy, textual fields. They assume uniform representation and cannot capture dependencies expressed through partial matches or semantically related phrases. In contrast, LLMs are well-suited for such cases, leveraging pretraining on large-scale natural language data to recognize subtle patterns and semantic signals in free-text attributes, enabling more effective tasks such as imputation.
III Preliminary
We now formalize the problem of missing data imputation, introduce the goal of LDI and provide an overview of our solution.
@@ III Preliminary
Problem Statement: Consider a dataset $\mathcal{D}$ with schema $\mathcal{R}$, consisting of attributes $A_{j}\in\mathcal{R}$. The dataset is represented as a collection of tuples, where each tuple $t\in\mathcal{D}$ contains a set of cells. A missing value in the $j$-th attribute of the $i$-th tuple is represented as $t_{i}[A_{j}]=\varnothing$. Let $\hat{\mathcal{D}}$ be an imputed version of $\mathcal{D}$, where for all $i,j$ such that $t_{i}[A_{j}]$ is missing in $\mathcal{D}$, there exists an imputed tuple $\hat{t_{i}}\in\mathcal{\hat{D}}$ where $\hat{t_{i}}[A_{j}]\neq\varnothing$. We denote $\mathcal{D}^{*}$ as the ground truth dataset, where each tuple $t_{i}^{*}\in\mathcal{D}^{*}$ contains the true cell values $t_{i}^{*}[A_{j}]$ for all attributes $A_{j}\in\mathcal{R}$.
@@
Goals: The goal of this paper is to take a dataset $\mathcal{D}$ and predict the entries in the target attribute $A_{T}$ for each tuple $t_{i}$ where $t_{i}[A_{T}]=\varnothing$, by focusing only on the most relevant attributes and tuples. We introduce a localized imputation method that uses the reasoning capabilities of LLMs while addressing key practical needs. Specifically, our method aims to (1) achieve high imputation accuracy in real-world datasets, even when the missing values depend on complex and unknown relationships, (2) improve explainability by making it possible to trace which parts of the input influenced the prediction and why they were selected, (3) handle common challenges such as formatting inconsistencies or noisy fields without relying on strict assumptions, and (4) minimize the need for external datasets or task-specific resources, allowing the method to be applied more broadly and efficiently.
Assumptions: Our method is based on the assumption that missing values can be inferred from repeated patterns and textual cues in the data. Specifically, we assume that (1) at least one other attribute is related to $A_{T}$ (the attribute with missing entries), meaning the missing values can be predicted from the wording, structure, or textual cues present in the other fields, and (2) some values of $A_{T}$ occur multiple times across different tuples, allowing the model to learn meaningful associations from this repetition. We discuss this further in the next section.
IV Proposed Method

### v2

@@ Abstract.
Missing values are pervasive in real-world tabular data and can significantly impair downstream analysis. Imputing them is especially challenging in text-rich tables, where dependencies are implicit, complex, and dispersed across long textual fields. Recent work has explored using Large Language Models (LLMs) for data imputation, yet existing approaches typically process entire tables or loosely related contexts, which can compromise accuracy, scalability, and explainability. We introduce LDI, a novel framework that leverages LLMs through localized reasoning—selecting a compact, contextually relevant subset of attributes and tuples for each missing value. This targeted selection reduces noise, improves scalability, and provides transparent attribution by revealing which data influenced each prediction. Through extensive experiments on real and synthetic datasets, we demonstrate that LDI consistently outperforms state-of-the-art imputation methods, achieving up to 8% higher accuracy with hosted LLMs and even greater gains with local models. The improved interpretability and robustness also make LDI well-suited for high-stakes data management applications.
Data Imputation, Dependency Evaluation, Explainability
1. Introduction
@@
Missing values are a common characteristic of many real-world datasets, particularly those derived from open or heterogeneous sources. If left unaddressed, they can significantly undermine the reliability of downstream analysis and machine learning applications. For example, many learning algorithms operate under the assumption that all input data is fully observed (Emmanuel et al., 2021; Zhang et al., 2008; Kumar et al., 2017), and their performance often degrades when faced with incomplete records. This not only reduces predictive accuracy but can also introduce systematic bias into analytical outcomes (Kang, 2013; Zhang et al., 2007; Qin et al., 2007; Stoyanovich et al., 2020). As a result, predicting or imputing missing values is a critical component of the broader data wrangling process, often helping to preserve the underlying structure of the data and ensuring the validity of subsequent analysis (Song and Sun, 2020).
Motivation. While the problem of data imputation has been studied for decades, a growing fraction of modern datasets are text-rich tables that mix structured fields with lengthy textual descriptions, reviews, or metadata. These tables are common in domains such as e-commerce, real estate, scientific data management, and open government data. Unlike purely numerical datasets, text-rich tables contain dependencies that are implicit, complex, and often buried within long, noisy text. Identifying which textual cues are informative for imputation remains a major open challenge.
@@
Example. Consider a restaurant table with missing values in the City column (see Figure 1). Several other attributes—including Description, Phone, and Reviews—might help infer the correct city, but their relevance varies. The Description column may contain location-related hints (e.g., “Modern American dishes” and “Elegant Italian dining”) that are typically too general or vague to be useful, while the Reviews field may offer more specific but noisy references (e.g., “Best gelato in Vegas”). The Phone column, in contrast, encodes reliable geographic signals via area codes, though inconsistencies in formatting can obscure them. Determining which attributes and records truly contribute to accurate imputation requires both contextual understanding and selective reasoning.
Figure 1. Example of data imputation
@@
Limitations of Existing Work. Existing approaches to data imputation range from simple heuristics to advanced generative models. Heuristic techniques (e.g., mean or mode imputation) are fast but ignore inter-attribute dependencies, often leading to biased estimates (Baraldi and Enders, 2010; Khan and Hoque, 2020). Model-based methods, including regression, expectation-maximization (EM), and matrix factorization, capture statistical dependencies but can be computationally expensive and rely on strong assumptions about the data distribution (Little and Rubin, 2019). Recent deep learning–based approaches, such as autoencoders, GANs, and transformer models, have shown promise in capturing nonlinear patterns, especially in large or multi-modal datasets (Yoon et al., 2018). Building on this progress, large language models (LLMs) have emerged as powerful tools for data imputation, thanks to their ability to model complex dependencies and generalize across diverse contexts (Narayan et al., 2022; Zhang et al., 2024; Li and Döhmen, 2024). However, current LLM-based approaches remain fundamentally global in scope as they tend to process entire tables or large portions thereof. This introduces three major issues that remain largely overlooked: (1) Scalability: Tables often exceed model context windows, forcing either truncation or splitting that loses important structure. (2) Noise sensitivity: Including loosely related rows and columns dilutes the signal, making imputations less accurate. (3) Lack of transparency: When the entire table is fed to an LLM, it becomes unclear which attributes or tuples influenced each prediction, undermining interpretability and user trust.
These limitations raise an important but underexplored question: Can missing values be accurately imputed using only a localized subset of the table—one that preserves essential context while improving scalability and explainability?
@@
Our Approach. We hypothesize that, in many cases, missing values can be imputed accurately using a localized context—a compact subset of attributes and tuples most relevant to the missing value. Our approach, LDI, operationalizes this idea by decomposing the problem into two sub-tasks: (1) selecting a subset of columns that are most relevant to the column with missing values, and (2) identifying a subset of rows that provide sufficient contextual evidence for imputation. This localized reasoning enables LDI to reduce noise, scale to larger tables, and trace which data influenced each prediction.
Our extensive evaluation across multiple datasets demonstrates that LDI not only improves performance but also enhances interpretability and robustness, particularly in text-rich, noisy, and heterogeneous environments.
Contributions: Our contributions can be summarized as follows: (1) We introduce LDI, a framework that performs attribute- and tuple-level selection to guide LLM-based imputation. (2) By limiting context to relevant subsets, LDI achieves both higher accuracy and transparent attribution. (3) LDI handles textual inconsistencies and heterogeneous formats more effectively than existing models. (4) The framework supports small, local LLMs, enabling efficient and privacy-preserving imputation. (5) Through analytical and experimental evaluations on real-world datasets from various domains, we demonstrate that LDI achieves superior accuracy, scalability, and explainability compared to state-of-the-art methods. (6) To support reproducibility and foster further research, we publicly release our code, framework, and datasets [^1] https://github.com/soroushomidvar/LDI.
2. Related Works
@@ 2.1. Data Imputation
Traditional Methods: Basic approaches replace missing values with fixed constants, such as zero or the mean, or use strategies like Last Observation Carried Forward (LOCF) and Next Observation Carried Backward (NOCB) (Engels and Diehr, 2003; Lachin, 2016; Klamroth-Marganska et al., 2014). These approaches are computationally inexpensive and straightforward to implement but rely on restrictive assumptions, such as the randomness of missing values or ordering of tuples, that often lead to biased results (Molnar et al., 2008), limiting their applicability in real-world scenarios where data is complex and irregular.
@@
Statistical Models: While basic methods rely on direct substitution, statistical models enhance imputation accuracy by estimating the underlying data distribution and predicting missing values accordingly. Examples include approaches based on relational dependency networks (Mayfield et al., 2010), factor graphs (Rekatsinas et al., 2017; Wu et al., 2020), distance likelihood maximization (Song and Sun, 2020), low-rank matrix completion (Candes and Recht, 2012), and non-linear models such as tree-based ensembles (Stekhoven and Bühlmann, 2012; Chen and Guestrin, 2016). While effective in certain scenarios, these approaches often rely on assumptions about data properties, such as normality or linearity, which can limit their applicability to real-world datasets.
@@
Similarity-based Approaches: Another line of research focuses on similarity-based techniques, where missing values are inferred based on similar neighbors with complete values. Neighbor-based methods, such as k-Nearest Neighbors (KNN) (Altman, 1992) and its extension, KNN Ensemble (KNNE) (Domeniconi and Yan, 2004), find the nearest neighbors with the smallest distance on complete attributes and aggregate their values as the imputation for the missing value. Similarly, MIBOS (Wu et al., 2012) imputes missing values through a similarity-based iterative approach that counts matching attribute values. Instead of relying on individual neighbors, some methods group similar tuples into clusters and perform the imputations within each cluster. For example, CMI (Zhang et al., 2008) imputes missing values by first clustering the dataset and then generating plausible values using similar tuples in the cluster. MGI (Yan et al., 2015) applies the Gaussian mixture model (GMM), and IFC (Nikfalazar et al., 2017) uses iterative fuzzy clustering to integrate mean imputation, hot deck, and multiple imputation. However, these similarity-based methods are constrained by their similarity functions, cannot handle semantic similarity, and have limited performance if used directly on the data without additional steps.
@@
Generative Models: Beyond these traditional methods, recent advances in generative models have opened new possibilities for data imputation by enhancing the ability to learn data patterns and generate more realistic values. For instance, GAIN (Yoon et al., 2018) uses the Generative Adversarial Nets (GAN) framework, where a generator imputes missing components and a discriminator ensures the generated data follows the true distribution. MIDA (Gondara and Wang, 2018) proposes a multiple imputation model using deep denoising autoencoders, and HI-VAE (Nazabal et al., 2020) adapts variational autoencoders to work with heterogeneous data. IPM (Mei et al., 2021) leverages pre-trained language models to capture semantics for predicting missing values in databases. It introduces IPM-Multi, which models imputation as multiclass classification, and IPM-Binary, which uses binary classification for candidate selection. LakeFill (Yang et al., 2025) extends the role of LLMs in imputation by incorporating external data lakes and focusing on tuple-level retrieval, addressing cases where local redundancy is insufficient. ZeroEC (Wu et al., 2025) focused on error correction using structured fields through LLM-based rule generation. UnIMP (Wang et al., 2025) introduces a framework for imputing mixed-type data by integrating LLMs with a message-passing network, and FMW (Narayan et al., 2022) also demonstrates that LLMs can handle missing data imputation effectively by framing them as prompting tasks. Although these models achieve state-of-the-art results, their black-box nature lacks explainability, making it difficult to understand how and why specific imputations are generated. The inability to trace an imputed value back to a specific part of the input data prevents validating, correcting, and refining imputations and limits the model’s applicability where understanding why a particular value was imputed can be just as important as the imputed value itself.
2.2. Dependency Detection
@@
In the context of attribute selection, our method is closely related to research on identifying dependencies between attributes, particularly through the use of Functional Dependencies (FDs) and Approximate Functional Dependencies (AFDs). FDs describe exact relationships where the value of one attribute is fully determined by another. Formally, an FD $X\rightarrow A$ holds in a relation $r$ if for all tuples $t$ and $u$ in $r$, $t[X]=u[X]$ implies $t[A]=u[A]$. FDs are widely used in database design to capture exact deterministic relationships (Papenbrock and Naumann, 2016; Bohannon et al., 2006; Salem and Abdo, 2016; Al-Janabi and Janicki, 2022; Ni et al., 2024). In practice, however, FDs are often too strict due to inconsistencies or noise in real-world data. To address this, AFDs have been introduced (Kivinen and Mannila, 1995; Mandros et al., 2017). An AFD $X\rightarrow A$ holds approximately if it is satisfied by most of the data, with the quality measured by the error $e(X\rightarrow A)$, defined as the minimum fraction of tuples that must be removed to make the dependency exact (King and Legendre, 2003).
@@
FDs and AFDs underpin many data cleaning techniques, including sparse regression for dependency discovery (Guo and Rekatsinas, 2019), scalable FD-based repairs for iterative data pipelines (Rezig et al., 2021), conditional FDs (Bohannon et al., 2006), and dynamic AFDs that tolerate dirty data by incorporating data source information (Ding et al., 2024). RENUVER (Breve et al., 2022) and Derand (Song et al., 2020) similarly leverage FDs to guide candidate selection for imputation. While effective on structured fields, these methods struggle with patterns embedded in free-text attributes and rely mainly on surface-level similarity rather semantic reasoning. They also assume consistent formatting and uniform representation, limiting their ability to capture dependencies expressed through partial matches or semantically related phrases. LLMs, in contrast, can model such dependencies more flexibly by drawing on their contextual understanding of language.
3. Preliminary
We formalize the problem of missing data imputation and outline the setting for LDI.
@@ 3. Preliminary
Problem Statement. Let $\mathcal{D}$ be a dataset with schema $\mathcal{R}$ consisting of attributes $A_{j}\in\mathcal{R}$. Each tuple $t_{i}\in\mathcal{D}$ contains a set of cell values, where a missing entry in the $j$-th attribute is denoted by $t_{i}[A_{j}]=\varnothing$. Let $\hat{\mathcal{D}}$ denote the imputed version of $\mathcal{D}$ such that, for all $i,j$ where $t_{i}[A_{j}]$ is missing in $\mathcal{D}$, the corresponding imputed tuple $\hat{t_{i}}\in\hat{\mathcal{D}}$ satisfies $\hat{t_{i}}[A_{j}]\neq\varnothing$. We denote by $\mathcal{D}^{*}$ the ground truth dataset in which each tuple $t_{i}^{*}\in\mathcal{D}^{*}$ contains the true values $t_{i}^{*}[A_{j}]$ for all attributes $A_{j}\in\mathcal{R}$.
@@
Goal. Given a dataset $\mathcal{D}$, the objective is to produce an imputed dataset $\hat{\mathcal{D}}$ by predicting missing entries of a target attribute $A_{T}$, reconstructing the ground truth dataset $\mathcal{D}^{*}$. LDI focuses on identifying the most relevant tuples and attributes for each imputation task and leveraging the reasoning capabilities of LLMs to achieve the following: (1) high imputation accuracy in real-world datasets, even when missing values depend on complex and unknown relationships; (2) improved interpretability by tracing which parts of the input influenced each prediction; (3) robustness to noise and formatting inconsistencies without relying on strict assumptions; and (4) minimal dependence on external data or task-specific resources, enabling broader and more efficient applicability.
Assumptions. We assume that (1) some values of $A_{T}$ (the attribute with missing entries) appear more than once across tuples, allowing the model to learn associations from these repetitions, and (2) one or more textual or categorical attributes are correlated with $A_{T}$ through textual, structural, or contextual cues. Similar assumptions are commonly made in prior work on data imputation (Rekatsinas et al., 2017; Breve et al., 2022; Rezig et al., 2021; Zhang et al., 2008).
4. Proposed Method